# RT-cache: Efficient Robot Trajectory Retrieval System

### v1

@@ IV-D Improved Pipeline Speed
Full-DB Search Time-out ($>$300 s)
@@
Figure 4: Scalability and example retrievals. Top: 3D plots compare a ground‐truth trajectory (black) with three retrieved motion snippets (colored lines) under different camera viewpoints, illustrating minor variations in length or shape. Bottom: Each row depicts a retrieved snippet from its initial “current” frame to subsequent steps, demonstrating how RT-cache aligns retrieval with each camera perspective (wrist, side, or front). Despite a growing memory (millions of states), retrieval remains sub‐linear in time thanks to hierarchical selection.
IV-E Discussion
IV-E1 How Does Retrieval Scale with Database Coverage?
The results in Table II show that multi-step retrieval effectively replaces costly repeated model calls and reduces task time. These findings underscore the importance of database coverage: having at least a few examples for each camera viewpoint and object arrangement is crucial to stable retrieval. Even a single new snippet can fix entire failure classes (e.g., side-camera views of the mug).
IV-E2 What Are Common Failure Modes, and How Can We Mitigate Them?
Figure 5 illustrates typical errors, such as minor positional offsets or incorrect gripper orientation. These near-misses often occur when an object is partially occluded or the snippet comes from a slightly different pose. Another challenge arises when the environment changes mid-execution (e.g., someone moves the bowl). Since RT-cache replays a snippet for up to $N$ steps, unmodeled scene shifts can cause collisions or overshoot. More frequent retrieval queries or auxiliary sensing (e.g., force feedback) could mitigate these issues.
IV-E3 How Similar Is the Retrieved Path to the Original?
  • Section A.1 (Three Trajectories). Figure 4 shows that each snippet (e.g., from the wrist-cam) targets the same object and viewpoint but can diverge slightly, much like a human “repeating” a skill with minor variations. This illustrates how our pipeline focuses on functionally correct paths rather than exact replicas.
  • Section A.2 (Similarity to Original). While some replayed paths are longer or shorter, they stay close enough to the ground‐truth trajectory (black line) to fulfill the intended goal (e.g., achieving a stable grasp).
  • Section B.1 (Next-Step Retrieval). The bottom images sequence each snippet from its “current” state onward, highlighting how RT-cache continuously selects forward-looking frames in line with the camera view.
@@ IV-E3 How Similar Is the Retrieved Path to the Original?
  • Section B.2 (Viewpoint‐Specific Retrieval). The system retrieves snippets aligned with the active camera: wrist‐cam queries yield wrist‐cam snippets, side‐cam queries yield side‐cam snippets, and so forth. This viewpoint consistency substantially improves retrieval accuracy in our trials.
IV-E4 How Does Horizon Length Balance Speed and Adaptation?
From Table III, short horizons ($N=1$) enable continuous correction but add overhead from repeated retrieval calls, whereas longer horizons ($N=5$) risk errors if a snippet is poorly matched. In many practical scenarios, $N=3$ provides a good trade-off. Future work could explore dynamic horizon selection, where the snippet length is automatically adjusted based on real-time confidence or scene changes.
Figure 5: Typical failure cases. (A) The arm halts just short of the mug’s graspable region. (B) The retrieved snippet steers the gripper too far right of the water bottle. (C) Another snippet overshoots left of the same bottle. (D) The gripper descends beyond the bowl’s rim, missing the target height.
@@ V LIMITATIONS AND FUTURE DIRECTIONS
Image-Only Retrieval.
While our approach matches image embeddings to retrieve multi-step actions, relying solely on visual similarity can be limiting. Future directions include incorporating additional feedback signals, such as human annotations via Reinforcement Learning from Human Feedback (RLHF) [28] or high-level textual cues (e.g., “grasp the bottle on the right”), to refine snippet selection. Integrating action requirements (e.g., language or state embeddings) could further disambiguate visually similar scenes. Moreover, adopting active perception strategies [29] and incorporating common-sense reasoning [16] may help the robot gather pertinent information, thereby improving retrieval quality for tasks that demand nuanced context.
Expanded Experimental Setup.
Although we focus on a single-arm manipulation scenario with a limited set of objects, broader evaluations are needed to fully gauge generalization and robustness. Future work will explore multi-object tasks, varying shapes and materials, and cluttered environments. Such tests will reveal the pipeline’s performance boundaries under more diverse conditions and guide domain-specific optimizations.
@@ VI CONCLUSIONS
We presented RT-Cache, a trajectory-memory pipeline that drastically reduces per-step inference costs by retrieving and reusing relevant trajectories from a large-scale robotic dataset. This approach embodies the principle of learning from experience, enabling tasks to be completed in a fraction of the time required by baseline Vision–Language–Action (VLA) models. While our experiments primarily focused on a grasp-like task, the pipeline is broadly applicable and can be integrated with standard humanoid controllers or higher-level planning frameworks. Ultimately, our aim is to bolster the real-world viability of big-data-driven robot systems, bridging the gap between the representational power of large VLA models and the stringent latency requirements of real-time operations.

### v3

@@ IV-A Performance Comparison
Full-DB Search Time-out (>>300 s)
@@
Figure 6: Single‑episode anchoring on a pushing task. Left: 3D end‑effector trajectories for ground truth vs. RT‑Cache. Colors denote motion segments; RT‑Cache closely tracks the reference path. Right: Third‑person snapshots with colored arrows indicating end‑effector motion. The memory initially contained no in‑domain data; adding one short episode (no training/fine‑tuning) created an anchor that RT‑Cache retrieves and replays, yielding a trajectory that qualitatively coincides with the ground truth. This illustrates retrieval‑as‑control: a single exemplar can unlock real‑time execution.
IV-B System efficiency and compute budget
Table I compares GPU/disk budgets and fine‑tuning needs. RT‑Cache runs in $\sim$8–10 GB GPU, shifts cost to storage ($\sim$100 GB embeddings+metadata in a vector DB), and requires no deployment‑time training or fine‑tuning. VINN and Behavior Retrieval train their own encoders/policies (BYOL; VAE+BC) with smaller footprints, while OpenVLA‑OFT is heaviest (16–24 GB GPU) and adds per‑checkpoint disk from fine‑tuning. Net effect: RT‑Cache trades SSD space for compute, amortizing model work into an append‑only memory that enables real‑time control on modest GPUs.
IV-C Ablation Studies (Fig. 5)
Object robustness
With a few in‑domain episodes, RT‑Cache is reliable across objects and views. Overall success is $96\%$ with the wrist‑cam setup (median time $66\,\mathrm{s}$) and $93\%$ from the third‑person front view ($63\,\mathrm{s}$), while the third‑person side view drops to $52\%$ ($77\,\mathrm{s}$). Per‑object totals span $78\text{--}100\%$ (bowl), $22\text{--}100\%$ (cup), and $56\text{--}100\%$ (bottle), with completion times between $60\text{--}68\,\mathrm{s}$ in the easier views. In contrast, the zero‑shot setting yields $0\%$ success across objects. Most errors are near‑misses rather than wrong‑turn failures (Fig. 5A).
Horizon sensitivity
In Fig. 5.B, varying the snippet length $N\in\{1,3,5\}$ trades reactivity for speed: N=1N{=}1 adapts every step but adds lookup overhead, N=5N{=}5 is fastest when the match is perfect yet less forgiving to small mismatches, and N=3N{=}3 consistently gives the best speed–accuracy compromise. Aggregate success/time values are in the caption. Unlike VINN (per‑timestep copying) or Behavior Retrieval (offline policy, no test‑time retrieval), RT‑Cache exploits N>1N{>}1 multi‑step replay without any training or fine‑tuning.
How similar is the retrieved path to the original in (Fig. 5.C)?
  • 1 — Object-Centric retrieval. Each retrieved snippet (e.g., wrist view) targets mostly same object and viewpoint but shows small variations—akin to a human repeating a skill with minor differences.
  • 2 — Temporal-Coherence retrieval. The bottom sequences illustrate progression from the “current” frame through subsequent steps, showing how RT‑Cache continually selects forward‑looking frames consistent with the active view.
@@ IV-B System efficiency and compute budget
  • 3 — Viewpoint‑specific retrieval. Queries respect the camera: wrist‑cam retrieves wrist‑cam snippets; side‑cam retrieves side‑cam, and so on. This viewpoint consistency noticeably improves retrieval accuracy in our trials.
Retrieval latency and throughput
Table II compares retrieval times for different database-selection strategies. An exhaustive nearest-neighbor search over millions of embeddings can take minutes, effectively making real-time tasks impossible. In practice, traditional systems might time out because they are not designed for massive datasets. Instead, our strategy completes lookups in under one second, even on large corpora, keeping RT-cache practical for real-world deployments where every second matters—such as human–robot collaborative settings.
Failure modes and mitigations
Figure 5.D illustrates typical errors, such as minor positional offsets or incorrect gripper orientation. These near-misses often occur when an object is partially occluded or the snippet comes from a slightly different pose.
Embedding keys (CLIP [23] vs. DINOv2+SIGLIP [21, 22])
In main experiments we use DINOv2+SigLIP concatenation to improve recall. Using image features as the retrieval key, DINOv2+SigLIP achieves roughly higher success than CLIP, especially from the wrist view, while CLIP struggles in third‑person‑front scenes.
IV-D Advanced task with single‑episode anchoring (Fig. 6)
Setup. We evaluate a more complex, contact‑rich tabletop pushing task from a third‑person view. With no initial in‑domain data, we insert one short episode by upserting its frame embeddings into the vector database and logging cartesian action deltas for replay.
Result. At test time, RT‑Cache matches the scene to that single exemplar and replays the $N$-step snippet end‑to‑end, without fine‑tuning or gradients, producing a trajectory that qualitatively coincides with ground truth. A single example is enough to run the behavior in real time.
@@ V LIMITATIONS AND FUTURE DIRECTIONS
RT‑Cache retrieves with image‑only keys. This keeps deployment simple and universal, but can yield near‑misses under occlusion, fine pose changes, or heavy clutter/dynamics. The controller also replays short $N$-step segments rather than composing longer, multi‑stage skills, and sustained growth of the memory raises storage/indexing efficiency questions.
We will broaden evaluation to multi‑object, cluttered and dynamic scenes, and report metrics that isolate retrieval quality from execution [24, 25, 26]. On the modeling side, we plan to add multimodal keys (proprioception/depth/language) with confidence‑aware [18, 20], dynamic horizons [19], and to pair retrieval‑first control with a lightweight VLA fallback or memory‑authoring policy. We will explore compression to keep the append‑only memory fast at scale.
@@ VI CONCLUSIONS
We introduced RT‑Cache, a training‑free retrieval‑as‑control pipeline that replays multi‑step snippets from a large vector memory; a hierarchical search keeps million‑scale lookups sub‑second, shifting cost from compute to storage and enabling real‑time control on modest GPUs. In tabletop reaching, RT‑Cache achieves higher success and $\sim$30% lower median completion time than VINN while avoiding cross‑domain failures seen with Behavior‑Retrieval–style policies; a moderate horizon (N=3N{=}3) balances speed and reactivity, and viewpoint‑aligned replay reduces drift. Using DINOv2+SigLIP keys improves recall at negligible runtime cost, and a single‑episode anchoring study shows immediate adaptation on a contact‑rich task with no fine‑tuning. The framework is readily extensible to multimodal keys and can interoperate with VLA policies as a fallback or for memory authoring.